### v2

# Lightweight Baselines for Medical Abstract Classification: DistilBERT with Cross-Entropy as a Strong Default

Abstract
The research evaluates lightweight medical abstract classification methods to establish their maximum performance capabilities under financial budget restrictions. On the public medical_abstracts corpus, we fine-tune BERT-base and DistilBERT with three objectives—cross-entropy (CE), class-weighted CE, and focal loss—under identical tokenization, sequence length, optimizer, and schedule. DistilBERT with plain CE gives the strongest raw (argmax) trade-off, while a post-hoc operating-point selection (validation-calibrated, class-wise thresholds) substantially improves deployed performance; under this tuned regime, focal benefits most. We report Accuracy, Macro-F1, and Weighted-F1, release evaluation artifacts, and include confusion analyses to clarify error structure. The practical takeaway is to start with a compact encoder and CE, then add lightweight calibration/thresholding when deployment requires higher macro balance.
I Introduction
Biomedical literature is expanding faster than manual curation can keep up, motivating automated systems for triage, classification, and summarization. While large language models perform well [1], their compute, latency, and privacy costs hinder deployment in healthcare settings governed by policies such as HIPAA [2]. We therefore emphasize strong, efficient, and reproducible baselines built on compact encoders such as DistilBERT [3].
Why compact baselines? Despite the excitement around large models, many real-world healthcare pipelines operate under strict cost-to-serve and governance requirements (on-prem, air-gapped or VPC-constrained deployments; auditable updates; predictable inference latencies). In such settings, compact encoders that are easy to fine-tune and calibrate often deliver a better accuracy–efficiency trade-off than heavyweight alternatives, particularly when the target task is well-posed and labeled at scale (e.g., abstract-level categorization). Our study therefore focuses on strong, reproducible baselines with minimal moving parts (two encoders $\times$ three losses, plus a lightweight post-hoc operating-point selection) to maximize clarity and reproducibility.
Relation to domain-adaptive pretraining. Domain-adaptive models typically improve in-domain text modeling: SciBERT targets scientific prose [4]; BioBERT and PubMedBERT leverage large biomedical corpora [5, 6]; ClinicalBERT adapts to clinical narratives [7]. Our question is complementary: how far can readily available compact encoders (BERT-base, DistilBERT) go on a public medical-abstract benchmark under identical, budgeted fine-tuning, and are common imbalance remedies (class weighting, focal loss) beneficial in this regime?
Dataset characteristics and imbalance. Abstracts mix background, methods, and outcomes, yielding fuzzy label boundaries even under moderate class skew. In such settings, reweighting or focal modulation may overemphasize ambiguous examples and harm macro balance; thus we treat plain cross-entropy (CE) as a strong raw default and empirically test alternatives.
Raw vs. deployed operating points. We report both raw (argmax) results and tuned results obtained via post-hoc operating-point selection: temperature scaling on validation [8] followed by class-wise threshold tuning, with thresholds frozen on test. Under raw decoding, CE is generally most stable; under tuned decoding, focal shows the largest gains—highlighting that training objectives and deployment thresholds should be evaluated separately.
Calibration and governance. Calibrated confidence is important for decision support; simple post-hoc calibration can mitigate miscalibration in pretrained Transformers without retraining [8], aligning modeling practice with auditability and privacy requirements [2].
Research questions. We study: (RQ1) which objective (CE, WCE, focal) best optimizes accuracy and Macro-F1; (RQ2) whether DistilBERT matches or surpasses BERT-base under identical budgets; (RQ3) how objectives shape per-class errors; (RQ4) sensitivity to practical knobs; (RQ5) efficiency gains (parameters, size, throughput); and (RQ6) which configuration performs best under a tuned deployment operating point.
II Related Work
Our work sits at the intersection of medical text classification, class imbalance, and model efficiency. The field has moved from feature-engineered models to fine-tuned Transformers tailored to scientific and biomedical text, including SciBERT, BioBERT, and ClinicalBERT [4, 5, 7]. New pretraining that links structured lab data with knowledge-guided learning is emerging [9]. Class imbalance is often addressed with resampling or cost-sensitive losses such as reweighting and focal loss [10, 11]. To reduce compute and latency, efficiency methods such as distillation (DistilBERT), pruning, and quantization are widely used [3]. Parallel efforts pursue interpretability and knowledge-enhanced agents for causal analysis in healthcare [12]. We compare compact, general-domain encoders with common imbalance objectives on a public benchmark to provide a practical baseline.
In a broader context, our work links to adjacent threads in medical AI. As models handle text, images, and other modalities, safety and auditing for multimodal systems are important; lightweight vision–language detectors for harmful content and targeted bias stress suites provide useful patterns [13, 14]. The question of model scale remains open, with mixed evidence on whether larger models yield better diagnostic value and on the trade-offs between performance, cost, and interpretability [15]. To manage growing system complexity, low-code, agentic frameworks can standardize pipelines and improve reproducibility [16]. For error analysis, beyond confusion matrices, topology-preserving projections from computational biology suggest helpful visualization tools [17].
III Methods
III-A Task and Dataset (HF medical_abstracts)
[…]
Splits.
If an official test split is present, we use it directly; otherwise, we stratify-split $10\%$ from the raw training portion to form a test set. From the training portion, we further hold out a stratified $10\%$ as validation using train_test_split [20]. Table I summarizes counts and per-class prevalence.
Preprocessing and tokenization.
We apply minimal, deployment-friendly preprocessing (strip empty lines/boilerplate and normalize whitespace). Tokenization relies on the model-matched uncased WordPiece tokenizer with right truncation and fixed-length padding to a maximum sequence length of 256 tokens for all runs. Padding uses padding='max_length' so every batch has uniform length. No stemming/lemmatization is used to preserve biomedical terms.
TABLE I: Dataset statistics for medical_abstracts. A stratified 10% validation split is held out from the official training set.
Train Validation Test
# Documents 10,395 1,155 2,888
# Classes 5
Per-class (names / %) Neoplasms Digestive system diseases Nervous system diseases Cardiovascular diseases General pathological conditions 21.91% 10.35% 13.33% 21.13% 33.28%
III-B Metrics
We report Accuracy, Macro-F1, and Weighted-F1 using scikit-learn implementations [21, 22]. In addition, we compute class-wise precision/recall/F1 and visualize error structure with confusion matrices and per-class bar charts [23, 24]. Unless otherwise noted, metrics are computed with the same evaluation code across all nine configurations on the held-out test split.
III-C Models and Objectives
We evaluate BERT-base-uncased and DistilBERT-base-uncased [25, 3] as encoder backbones, each followed by a randomly initialized linear classification head (hidden size 768, output size 5). The encoders are fine-tuned end-to-end.
Loss functions.
Let $p_{t}$ be the predicted probability of the true class $t$. We compare:
$\mathcal{L}_{\mathrm{CE}}$ $=-\log p_{t},$ (1)
$\mathcal{L}_{\mathrm{WCE}}$ $=-\,w_{t}\log p_{t},$ (2)
$\mathcal{L}_{\mathrm{FL}}$ $=-\,\alpha_{t}\big(1-p_{t}\big)^{\gamma}\log p_{t},$ (3)
with $\gamma{=}2.0$ for focal loss [10]. For weighted cross-entropy (WCE), class weights $w_{t}$ are derived from training-split frequencies via an inverse-frequency scheme (normalized). For focal loss (FL), we use class-dependent $\alpha_{t}$ computed from inverse class frequency (normalized to unit mean). All other settings (tokenizer, max length, optimizer budget) are kept identical across objectives to isolate the effect of the loss.
III-D Training and Implementation Details
We fine-tune with AdamW [26] using a learning rate of $2{\times}10^{-5}$, batch size $16$, and $3$ epochs per run. We employ the default linear learning-rate schedule with 500 warmup steps. Evaluation is performed at the end of each epoch on the validation split, and the best checkpoint (by validation Macro-F1) is automatically restored at the end of training. Unless otherwise specified, we use a single fixed seed for each run. Implementation relies on Transformers and Datasets [27, 28].
Outputs and artifacts.
For each raw configuration (BERT/DistilBERT $\times$ {CE, WCE, FL}) and each tuned DistilBERT configuration (CE/WCE/FL), we export per-run JSON/CSV metrics (including global and per-class scores), confusion matrices, per-class metric plots, and model/tokenizer checkpoints, together with the selected thresholds for the tuned runs. These artifacts enable exact reproduction of all numbers reported in Section IV.
Operating Point Optimization (post-hoc)
After training, we optionally fit temperature scaling on the validation split [29] and perform a small grid search of class-wise decision thresholds to maximize validation Macro-F1. The selected temperature (often T=1T{=}1 when unhelpful) and thresholds are frozen and applied once to the test set (no refitting). We therefore report both Raw (argmax) and Tuned (class-wise thresholds, with temperature applied only when beneficial) results to decouple training quality from deployment-time operating-point choice.
IV Experiments and Results
IV-A Experimental Setup
We evaluate six raw configurations formed by two encoders (BERT-base and DistilBERT) crossed with three objectives: cross-entropy (CE), class-weighted CE (WCE), and focal loss (FL). In addition, we report three tuned operating points for DistilBERT (CE/WCE/FL) obtained by post-hoc class-wise thresholding selected on validation and frozen on test. Unless noted, we fine-tune with AdamW (learning rate $2{\times}10^{-5}$), batch size $16$, for $3$ epochs; tokenization is uncased with truncation/padding at $256$ tokens. We adopt the official train/test split of medical_abstracts and reserve $10\%$ of the training set as a stratified validation split. All random seeds are fixed; we save the best checkpoint by validation Macro-F1 and evaluate on the test set. Metrics and plots are computed with scikit-learn utilities [21, 22, 23, 24]. The code relies on Transformers/Datasets [27, 28].
From a design perspective, this setup isolates the contribution of the objective under a constant optimization budget and identical tokenization policy. Using a stratified validation split reduces the variance of model selection when labels are unevenly distributed [30], and reporting Macro-F1 alongside accuracy is standard practice for skewed text classification [31, 32]. We include WCE and FL because they are widely used for class imbalance in deep classifiers [31, 11], while keeping training time comparable across losses to ensure a fair comparison.
IV-B Preprocessing and Tokenization
We keep the dataset text as-is (no lowercasing beyond the uncased tokenizer), retain punctuation, and strip empty lines. Sentences longer than $256$ tokens are truncated on the right; short sequences are padded to match the batch. Tokenization artifacts, special tokens, and attention masks follow library defaults [27]. Because abstracts often contain multi-part rhetorical structure (background, methods, results), a moderate maximum length ($256$) offers a good balance between coverage and speed. The fixed-length policy also stabilizes training dynamics by avoiding variability in effective batch compute, which is useful when comparing objectives head-to-head under the same budget.
IV-C Aggregate Performance
Table II summarizes results. Raw/argmax: DistilBERT with standard CE achieves the best balance (Accuracy $64.61\%$, Macro-F1 $64.38\%$, Weighted-F1 $63.25\%$), narrowly surpassing BERT-base with CE. Tuned (val-selected thresholds, frozen on test): focal loss benefits most from operating-point selection and attains the highest deployed scores (Acc. $\geq$ 80%), whereas CE also improves but to a lesser degree. (see Obs. O4) Figure 1 visualizes the Macro-F1 ranking across all nine settings. Figure 2 contrasts parameter count against Macro-F1 and shows that DistilBERT—the compact encoder distilled from BERT [3, 25]—occupies a favorable accuracy–efficiency region.
Observation O1 (Objective choice). Under raw/argmax decoding, WCE and FL underperform CE on this corpus, consistent with balanced-to-moderately-imbalanced label priors [18]. This aligns with broader evidence that when skew is mild and label ambiguity is present, cost-sensitive reweighting or focusing terms can over-amplify noisy/ambiguous instances and hurt precision [31, 11].
Observation O2 (Encoder choice). DistilBERT matches or slightly surpasses BERT-base at substantially lower capacity, supporting compact baselines as strong defaults when compute or latency is constrained [3].
Observation O3 (Stability). Across runs, the ranking (DistilBERT+CE $\succ$ BERT+CE $\succ$ {WCE, FL}) is consistent with the per-class evidence in Figure 4 and with prior findings that objective-driven gains should be verified against label noise and class prevalence before adoption [31].
Observation O4 (Decode mode matters). Objective rankings differ between raw and tuned settings: CE is the most stable raw default, while focal gains the most from post-hoc thresholds and yields the best deployed performance.
IV-D Per-class Behavior and Error Structure
The $\times3$ confusion panel (Figure 3) reveals strong diagonals for Class_1 and Class_4, while Class_5 shows distributed off-diagonal mass. The per-class bars (Figure 4) expose how objectives shift precision/recall:
  • Stable classes. Class_1/Class_4 remain robust across losses and encoders, likely due to distinctive lexical cues common in the abstracts of those categories.
  • Fragile class. Class_5 suffers recall deficits and spillovers into Class_4; many errors correspond to abstracts mixing background with outcome statements, which blurs boundaries.
  • Redistribution rather than reduction. WCE/FL mildly reshuffle errors across neighboring classes but rarely reduce global error mass, explaining their weaker Macro-F1 compared with CE. This pattern is consistent with reports that naive cost-weighting may shift, rather than shrink, confusions when minority classes are not cleanly separable [11].
IV-E Loss Comparison: Why Does CE Win Here?
Given the true-class probability $p_{t}$, all three objectives reduce to scaled variants of $\log p_{t}$:
$\mathcal{L}_{\mathrm{CE}}$ $=-\log p_{t},\quad\mathcal{L}_{\mathrm{WCE}}=-\,w_{t}\log p_{t},$ (4)
$\mathcal{L}_{\mathrm{FL}}$ $=-\,\alpha_{t}(1-p_{t})^{\gamma}\log p_{t},\ \ \gamma{=}2.0.$ (5)
On medical_abstracts, (i) skew is moderate rather than extreme [18]; (ii) many “hard” examples are ambiguous, not merely rare. Overweighting such points (via $w_{t}$ or $(1-p_{t})^{\gamma}$) can amplify label noise and degrade precision, explaining CE’s advantage [31]. While more sophisticated imbalance objectives exist (e.g., effective-number class-balanced losses) [33], our results indicate that, for this corpus and budget, a plain CE baseline is a strong and reliable choice.
Training dynamics, sensitivity, and calibration. Validation Macro-F1 improves mainly in the first two epochs, with stable AdamW optimization and no divergence at our settings, consistent with short-budget fine-tuning of pretrained Transformers [25]. We use the same evaluation code across runs [27]; model selection on a stratified validation split reduces variance [30]. A compact sweep with three seeds and small grids over sequence length $\{128,256,512\}$ and learning rate $\{2,3,5\}\times 10^{-5}$ is sufficient to confirm ranking and guard against configuration luck [31, 32]. For deployment, we report Expected Calibration Error and apply post-hoc temperature scaling on a held-out split to improve confidence calibration without retraining [29].
IV-F Efficiency and Footprint
DistilBERT’s distilled architecture [3] reduces parameter count to $\sim 66$M from BERT-base’s $\sim 110$[25], lowering memory footprint and cold-start latency. Figure 2 visualizes the accuracy–efficiency frontier on this task: the compact model attains equal or better Macro-F1 with smaller on-disk checkpoints, which simplifies on-prem deployment and resource isolation under privacy regulations such as HIPAA [2].
IV-G Takeaways for Practitioners
  • Start simple: DistilBERT+CE is a robust, compute-efficient baseline; treat WCE/FL as hypotheses to test, not defaults.
  • Inspect class behavior: use Figure 4 and Figure 3 to target fragile classes rather than inflating loss weights globally.
  • Mind calibration: apply temperature scaling and report ECE/Brier for decision-support scenarios [29].
  • Track cost-to-serve: pair accuracy with footprint/latency considerations (Figure 2) for deployability under privacy constraints [2].
TABLE II: Overall results on medical_abstracts: Accuracy, Macro-F1, and Weighted-F1 (%). We report raw (argmax) and tuned (validation-selected class-wise thresholds, frozen on test) results.
Configuration Acc. Macro-F1 Wtd-F1 Raw baselines (argmax) bert-base-uncased cross_entropy 64.51 63.85 62.12 bert-base-uncased class_weight 62.88 62.43 59.66 bert-base-uncased focal 62.81 62.47 59.93 distilbert-base-uncased cross_entropy 64.61 64.38 63.25 distilbert-base-uncased class_weight 62.29 62.22 59.24 distilbert-base-uncased focal 62.64 62.40 59.74 Post-hoc threshold-tuned (thresholds selected on val, fixed on test, optimized parameter 10e, 512, 3e-5, bs16) distilbert-base-uncased CE + tuned (all thresholds = 0.7) 74.03 70.73 72.42 distilbert-base-uncased Focal + tuned (all thresholds = 0.7) 82.22 77.55 81.22 distilbert-base-uncased ClassWeight + tuned (all thresholds = 0.7) 68.60 66.06 62.97
Figure 1: Macro-F1 for all settings. Left (gray): six raw configurations (BERT/DistilBERT $\times$ {CE, WCE, FL}). Right (blue): tuned configurations with class-wise thresholding (selected on validation and frozen on test). Tuned DistilBERT+Focal attains the best Macro-F1 (77.55%), tuned CE reaches 70.73%. See Table II for exact values.
Figure 2: Parameter count (millions) vs. Macro-F1. Color encodes loss (CE, Focal, Class Weight); marker shape distinguishes regimes (Raw, Tuned via per-class thresholding on val); fill indicates model (hollow=BERT $\approx$110M, solid=DistilBERT $\approx$66M). Tuned points lift Macro-F1 without changing parameter count.
Figure 3: Nine confusion matrices in a $\times3$ grid: top—BERT, middle—DistilBERT, bottom—DistilBERT (tuned). Columns: CE / WCE / FL.
Figure 4: Per-class precision/recall/F1 across configurations.
V Discussion, Limitations, and Impact
We decouple the training objective from the deployment operating point. Under moderate class skew and label ambiguity, upweighting hard/minority examples can amplify noise and reduce precision; accordingly, with raw/argmax decoding that isolates loss effects, cross-entropy (CE) remains the most stable and strongest on this corpus [31, 34, 11]. Focal loss is useful in long-tail vision settings [10], but its $(1-p_{t})^{\gamma}$ term may over-emphasize uncertain points when ambiguity—not rarity—is dominant. While alternatives like class-balanced losses or logit adjustment can help in different regimes [33, 35], the label distribution of medical_abstracts and our fixed budget make CE the most reliable raw choice [18]. Why then does focal score highest after tuning? The large gains arise from post-hoc operating-point selection (class-wise thresholds selected on validation and frozen on test), a generic step that can lift multiple models; this reflects improved decision thresholding rather than necessarily better pre-softmax representations. Given CE’s hyperparameter-free simplicity, robustness, and auditability, we retain it as the strong raw default, and recommend lightweight calibration/thresholding at deployment time when Macro-F1 is prioritized.
This study has limits. We use one public corpus and practical defaults; more tuning could change absolute numbers. A stratified validation split reduces selection variance but does not remove it [30]. Results may not transfer to clinical notes or radiology reports due to domain shift, a known risk in medical AI [36, 37]. We focus on head metrics and confusion patterns; we do not report subgroup fairness or prospective robustness, and calibration is addressed only with post-hoc scaling, so further checks are needed before clinical use, consistent with reporting guidance [38]. On impact, compact encoders lower cost and latency and fit on-prem deployments with privacy constraints such as HIPAA [2], but any use in care settings should remain decision support within clinical governance frameworks, with ongoing monitoring for shortcut biases and distribution shift [38, 36].
VI Conclusion and Future Work
We show that a simple, compute-efficient baseline, DistilBERT with cross-entropy, matches or slightly exceeds BERT-base at lower cost. On the medical_abstracts corpus [18, 19], class-weighting and focal loss do not surpass CE under identical budgets with raw/argmax decoding; with post-hoc operating-point selection, focal yields the highest deployed scores, highlighting the practical value of threshold selection. A practical path is to begin with a compact encoder and cross-entropy, then add calibration and explicit checks for dataset shift and governance before moving to heavier models [37, 38].
Future work includes extending to multimodal inputs from figures and tables while building strong safety and bias audits, informed by lightweight vision–language detectors and targeted bias stress suites [13, 14]. It is also useful to test compact encoders against larger systems reviewed in recent surveys and to examine scale versus clinical utility [39, 15], alongside efficient learning methods that lower parameters and cost [40]. To improve the workflow, agentic low-code frameworks can standardize pipelines and increase reproducibility [16], and topology-preserving projections can enrich error analysis beyond confusion matrices [17]. Architectural and optimization changes such as Bayesian-optimized attentive networks are another direction [41]. Finally, moving beyond static abstracts to longitudinal prediction [42] and exploring federated learning with data harmonization for privacy and non-IID settings remain important [43]. A systematic study of operating-point selection across objectives and datasets—beyond temperature and simple threshold sweeps—remains an important direction for robust deployment.